OpenStudy (anonymous):

Solve x2 – 4x = 12 by completing the square

mathslover (mathslover):

\[\large{x^2-4x-12=0}\] right?

OpenStudy (anonymous):

you just add 4 to both sides. that will make it \[(x-2)^{2}=16\]

OpenStudy (anonymous):

@mathslover yes

mathslover (mathslover):

ok so now .. \[\large{x^2-4x-12=x^2-4x+4-16=0}\] right?

OpenStudy (anonymous):

Yeah

mathslover (mathslover):

now can u write \[\large{x^2-4x+4=(x-2)^2}\] ??

mathslover (mathslover):

(x)^2 -2(2)(x)+2^2 = a^2-2ab+b^2=(a-b)^2=(x-2)^2

OpenStudy (anonymous):

Stuck

mathslover (mathslover):

ok so : (x^2-4x+4) = (x-2)^2 ?

OpenStudy (anonymous):

x^2-4x-12=0 x=-2 x=6? or am I doing it wrong

mathslover (mathslover):

I think you are doing it wrong ; : x^2-4x-12 = 0 \[\large{x^2-4x+4-16=0}\] \[\large{(x-2)^2-16=0}\] \[\large{(x-2-4)(x-2+4)=0}\] \[\large{(x-6)(x+2)=0}\] hence either x = 6 or x = -2

mathslover (mathslover):

always remember that : \[\large{(a-b)^2=a^2-2ab+b^2}\] \[\large{(a+b)^2=a^2+2ab+b^2}\]

mathslover (mathslover):

or : \[\large{(a\pm b)^2=a^2\pm 2ab +b^2}\]

OpenStudy (anonymous):

You're awesome! Thank you so much! The formulas were very helpful :)

mathslover (mathslover):

nice to hear .. :) best of luck